Actions
Ops/Comm Scheduling meeting : 28/2/2017¶
Location : CR21 in R1 at 2pm
Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402
- SS Magnet status
- Hydrogen status
- Operations
- MOM report
- Water chiller status
- Diffuser status
- Do we remove the diffuser and try to fix it in the between the user cycles, or just fix iris 4 open in situ for ever?
- We have 10 weeks from end of running on March 6 to start of next user cycle on May 2
- From John : Fix in Oxford : 3 to 6 weeks depending on what's wrong
- To fix iris 4 open : 1 day to 1 week
- Do we remove the diffuser and try to fix it in the between the user cycles, or just fix iris 4 open in situ for ever?
- FCD vacuum gauge readings
- Running plans to March 6th (end of running)
AOB:
- Issues List
- Last meeting : Tuesday Feb 21 2017
- Next meeting : Tuesday Mar 14 2017
Minutes¶
- CC Magnet status
SSD steady at 206 A; SSU is presently ramping (107A); will then ramp FC to 106 A, lower current than before, but higher SSU M2.
Josef Boehm is concerned that leakage in SSD may depend on polarity. Need to schedule a check of this, and a decision of which mode for LH2 running.
- Hydrogen status
Working on the mods for the insert, and waiting on getting the Hall next week. Start by taking out the loading platform and FC. Someone should make a crude check that LiH hasn't moved. Chris Rogers would like to see FC geometry for himself - plan in advance, no tourism!
The paperwork continues to grow; there will be a review on 20 March which should sign off on the proposed test regime, operating procedures, etc. Still then need to carry out said tests...
- Operations
- MOM report - see slides
EMR board swap has cured the DAQ failures. Now have no spare EMR board; this is a custom item put together by INFN.
MU has changed parts on some of the Tracker Hall probes, Tracker 4 now works but the other two (1 on each SS) still broken.
- MOM report - see slides
- Water chiller status
A mains spike (which caused damage around site last week) upset chiller 1 such that it refuses to talk to the Manager, so we no longer have auto-switchover.
Have moved the conventional magnets on to the loading-bay chiller.
The CC magnet cryocompressors are a constant load, reducing the need for the roof chillers to keep starting/stopping.
If the roof water fails, we can change the CC magnet cryocompressors back to the loading-bay chiller to keep the magnets cold, at cost of not being able to run (no conventional magnets ). Still need the change-over document (John Govans)
The Trench water incident - this will be reported to SHE [has been - I04731]. Should do a cut-out switch test on all Sumitomos after the run, also buying a couple of spares (at present have one borrowed spare - i.e. unused - in Hall). An expert group (Craig/Colin/John G ?) should investigate further.
- Diffuser iris 4:
Steady stream of visitors have inspected it today: yes, it doesn't open fully. Possibly a screw trying to escape from one of the brass petals.
V Blackmore will disconnect air supply from iris 4 to ensure it stays closed until the end of the run.
- FCD vacuum gauge readings
Insulating vacuum spikes from 10^-8 to 10^-6 and back - change is too quick to be real.
- Running plans to March 6th (end of running)
AOB:
- The standing "Detector Performance" item is missing.
There's a "hole" in the Tracker. Can we cure this by moving waveguides around? (M Uchida)
- There might still be money for spares, etc.
- Need to start preparing the shut-down work list (Colin Whyte), which in turn depends on the encountered issues list from Operations (Steve Boyd)
Actions¶
- JBoehm to get SS operating info from PHanlet (ongoing)
- JBoehm to get/impose consensus from SC Magnets group on running of "Central CC GUI"
- JBoehm arrange tests of polarity effect on SSD leaks
- Hall Manager when removing FC to make a crude check that LiH hasn't moved.
- Hall Manager: water system change-over procedure
- Status of
miceraid5
disk cleanup (stands) - DC to ensure Health Physics have checked temporary neutron monitor (stands)
- SBoyd - ensure list of issues arising from current running is being collated
- CWhyte - draw up the shutdown task list
- SBoyd - ensure MICE-wide discussion of MOM C&M panel
- PFranchini - Get all shifters to go through the CE checklist during Control Room refresher training
- MUchida - investigate moving Tracker waveguides around to cure "hole"
VBlackmore - disable diffuser iris 4(done)
Updated by Nebrensky, Henry over 6 years ago ยท 20 revisions